OpenStudy (anonymous):

the value pi/12 is a solution for the equation 2 cos^2(4x)-1=0

OpenStudy (anonymous):

\[2\cos ^24(\frac{ \pi }{ 12 })-1=0\]

OpenStudy (anonymous):

VERIFY LHS=RHS